OpenStudy (anonymous):

Need to find if these are true or false. I miss one here and there and can't figure out which. 1.The composition of an even and an odd function is even 2. The ratio of two odd functions is odd 3. The product of two even function is even 4. The composition of an odd function and an odd function is even 5. A function cannot be both even and odd. 6. The sum of two even functions is even 7. The sum of an even and an odd function is usually neither even or odd, but it may be even. 8. The product of two odd function is odd

OpenStudy (amistre64):

which ones are you missing out on?

OpenStudy (anonymous):

That's the thing, I don't know for sure, since my homework will only tell me if I have everything correct, not if I have 1 right here and there. I'll list the answers I have. T T T T F T F F

OpenStudy (amistre64):

6 and 7 might need work the rest seem fine working with the example: f = x, is odd g = x^2, is even

OpenStudy (amistre64):

5 and 7 that is

OpenStudy (amistre64):

f(x) = 0 is even and odd, its symmetric about the origin, and the y axis

OpenStudy (anonymous):

That's I figured that one would be true, since 0 is both even and odd. There was even a problem in the homework asking if there is a function that is both even and odd. So I have True for that one.

OpenStudy (amistre64):

f(x) = 0, f(-x) = 0 so f(x) = f(-x), its even f(-x) = 0, -f(x) = 0 so f(-x) = -f(x), its odd

OpenStudy (amistre64):

7 is just too wordy for me

OpenStudy (anonymous):

So that one seems to false, right, since it is saying that it cannot be both. That's the answer I had

OpenStudy (amistre64):

ah yes, wording counts lol 5. A function cannot be both even and odd. -------- id pick false

OpenStudy (anonymous):

Ya, so if everything looks good except the two you mentioned, I tried reversing number 7 but it is saying I still have one wrong.

OpenStudy (amistre64):

1.The composition of an even and an odd function is even f(g) = (x)^2 is even, T 2. The ratio of two odd functions is odd f/g = x/x^2 = 1/x, is odd g/f = x^2/x = x is odd, T 3. The product of two even function is even gg= x^2(x^2) = x^4, T 4. The composition of an odd function and an odd function is even f(f) = (x) is odd, so F might be this one

OpenStudy (amistre64):

lets work 7 a little let f be odd and g be even, and assume their sum can be even h(x) = f(x)+g(x) h(-x) = f(-x) + g(-x) = -f(x) + g(x) this is only even if f(-x) = f(x), but f is odd, and therefore is false by contradiction

OpenStudy (anonymous):

So, with your help we get T T T F F T F F ?

OpenStudy (anonymous):

When I do that it is still wrong =/. I had these type that don't give me a feel for what I did wrong or right. I know most of these, but now I'm questioning everything!

OpenStudy (amistre64):

id go with the last one, but if its saying wrong then im not sure where it would be faulty at

OpenStudy (amistre64):

it would take to long to proof them out, since im losing internet soon

OpenStudy (amistre64):

change 7 to T, just as a guess ... but im at a loss for it

OpenStudy (anonymous):

All right, thanks for trying to help any way. Got me on a good track!

OpenStudy (amistre64):

good luck with it ;)